Which of the following values represent the definite "{\int_{\pi/2}^{2\pi/3}\frac{1}{1+\cos{\theta}}d\theta}"



1) e


2) 1


3) "\sqrt{3}"


4) "\sqrt{3}-1"


1
Expert's answer
2021-08-23T16:33:09-0400
"\displaystyle\int_{\pi/2}^{2\pi/3}\dfrac{1}{1+\cos \theta}d\theta=\displaystyle\int_{\pi/2}^{2\pi/3}\dfrac{1}{2\cos^2(\dfrac{\theta}{2})}d\theta"


"=[\tan(\dfrac{\theta}{2})]\begin{matrix}
 2\pi/3 \\
\pi/2
\end{matrix}=\tan(\pi/3)-\tan(\pi/4)=\sqrt{3}-1"

4) "\sqrt{3}-1"